# Scaling Efficient Masked Image Modeling on Large Remote Sensing Dataset

### v4

@@ Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-13M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-13M comprises 13 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments demonstrate that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models.
@@ 1 Introduction
Figure 1: Advantages of SelectiveMAE on finetuning performance and pre-training efficiency. The accuracy is evaluated on the DIOR detection dataset [41], while the training efficiency is measured on NVIDIA A100 GPUs, with different versions of ViT [19].
Over the past decade, advancements in remote sensing (RS) technology and data acquisition have improved applications in ecosystem monitoring [54], natural disaster management [6], among others [32, 74]. These applications rely on essential capabilities such as scene classification [30, 14], object detection [43], change detection [85], and semantic segmentation [93]. However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
Figure 2: Advantages of OpticalRS-13M in data volumn and diversity. (a) The volumn of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets. (b) OpticalRS-13M includes significantly more sample categories than others, spanning both object and event types. (c) OpticalRS-13M provides more abundant feature representations compared to other large-scale RS datasets. Here, 1,000 samples are separately selected from different datasets, and features are extracted using a pre-trained ViT-B.
Thanks to advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [28, 89], the pre-training of visual foundation models has seen remarkable improvements [90, 94, 91, 95, 73, 77]. Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [70]. However, two challenges persist in the development of RSFMs. (i) Compared to the ImageNet-21k [58] dataset, previous RS datasets [48, 50, 49, 62] contain fewer samples (approximately 1 million vs. 14 million), which impedes the sufficient MIM training of large backbones. (ii) These datasets, which primarily focus on global scene semantics [48, 50, 49], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks. This limitation restricts the generalization of the learned representations.
@@
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [70, 65]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches (e.g., $\leq$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens (e.g., $\leq$50%) with higher HOG values for feature encoding (e.g., $\leq$15%) and pixel reconstruction (e.g., $\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence (Fig. 1). Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
@@
1) We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
2) Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
3) We introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training. It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
Experiment results suggest the effectiveness and scalability of the proposed pipeline. OpticalRS-13M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [28].
@@ 2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [62, 84], object detection [16, 21, 38], and segmentation [2, 76, 68]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. The SEN12MS dataset [59] consists of 180,662 triplets, each containing synthetic aperture radar (SAR) and multispectral Sentinel-2 image patches, and MODIS land cover maps. While SSL4EO-S12 [79] , SSL4EO-L [61] and SatlasPretrain [5] contain millions of images, the major data type in these datasets are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [80] and RS5M [96], but they primarily focus on multimodal tasks. Currently, there are also some datasets that focus solely on visible light RS images. MillionAID [48], SeCo [49] and CACo [50] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, with the proposed pipeline, we introduce the OpticalRS-13M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
@@
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [69]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [1, 50, 49, 81], recent advancements have primarily centered around generative self-supervised methods, especially in MIM. Specifically, many studies aim to improve generative self-supervised algorithms by leveraging general image knowledge [51, 33, 65], scaling up parameter sizes [10, 18], integrating spatio-temporal information [17, 92, 42], encompassing geometric attributes [70, 46], handling multi-sensor data [34, 24, 27, 8, 31], and employing multi-scale concepts [56, 64, 53]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
@@ 3.1 Dataset Curation
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this, we curate a large-scale RS dataset, OpticalRS-13M, through extensive data collection and preprocessing.
Data Collection We began by reviewing publicly available RS datasets from the past decade. Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [48]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity. Intra-class Diversity: High within-class diversity ensures comprehensive representation of real-world target distributions. We include data from different sources of the same scene or target, incorporating change detection data to capture object variations by geographic location and imaging time. Inter-class Diversity: To enable self-supervised pretraining to distinguish categories effectively, we included as many fine-grained categories as possible, especially those with high semantic overlap, helping the model learn robust feature representations. 2) Richness: Beyond diversity, richness is essential. We ensure varied content features and large sample sizes by collecting images under multiple conditions (e.g., weather, seasons, lighting), introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion. Sampling across imaging conditions provides a comprehensive real-world representation, enhancing model’s representation and generalization capabilities. 3) Scalability: Scalability is crucial given the evolving applications of remote sensing images. We continuously expand the dataset based on scene and target categories across classification, detection, and segmentation tasks. In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios.
Data Preprocessing Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained. To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
1) We focused exclusively on visible light images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training. 2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices. 3) We filtered out images with fewer than 64 or more than 1024 pixels. 4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [55], and then a refined phase involving manual review. This ensured that only highly similar images were excluded. As a result of these steps, we created the OpticalRS-13M dataset, which contains approximately 13 million high-quality RS images, significantly larger than previous representative RS datasets, as illustrated in Fig. 2 left. In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality. We’ve screened and processed most small and medium-scale datasets from the past decade in this field.
@@
Dataset Diversity To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig. 2 right. The analysis reveals that OpticalRS-13M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical remote sensing datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and so on. This expanded information enhances the versatility of OpticalRS-13M for a wider range of downstream tasks.
@@ 3.2.1 Preliminaries of MAE
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder. The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
@@ 3.2.3 Progressive Semantic Token Selection
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Inspired by the curriculum learning  [78, 7, 26], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig. 3. In this module, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. We define the initial set of selected token as:
@@
$\small\mathcal{D}(S^{U},S^{I})=\mathbf{1}-\cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-0.96873pt\left|\mathbb{S^{U}}\right|\kern-0.96873pt\right|\cdot\left|\kern-0.96873pt\left|\mathbb{S^{I}}\right|\kern-0.96873pt\right|},$ (3)
@@
Figure 5: Illustration of the patch selection process in PSTS.

### v5

# Harnessing Massive Satellite Imagery with Efficient Masked Image Modeling

@@ Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-13M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-13M comprises 13 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments show that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models. The dataset, source code, and trained models will be released at SelectiveMAE.
@@ 1 Introduction
Figure 1: Advantages of SelectiveMAE on finetuning performance and pre-training efficiency. The accuracy is evaluated on the DIOR detection dataset [1] with different versions of ViT [2]. The training efficiency is measured by Data Throughput/Minute, i.e., the processed images per minute, on 8 $\times$×NVIDIA A100 GPUs.
Over the past decade, advances in remote sensing (RS) technology and data acquisition have enhanced applications in ecosystem monitoring [3], natural disaster management [4], among others [5, 6]. These rely on essential capabilities such as scene classification [7, 8], object detection [9], change detection [10], and semantic segmentation [11]. However, each task typically demands substantial computational resources to train specialized models and learn task-specific feature representations.
Figure 2: Advantages of OpticalRS-13M in data volumn and diversity. (a) The volume of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets. (b) OpticalRS-13M includes significantly more sample categories than others, spanning both object and event types. y-axis denotes the data volume in (a-b). (c) OpticalRS-13M provides more diverse feature patterns than other large-scale RS datasets. Here, 1,000 samples are separately selected from different datasets, and their features from a pre-trained ViT-B are visualized using t-SNE [12].
Recent advances in self-supervised learning, particularly Masked Image Modeling (MIM) techniques [13, 14], have significantly improved the pre-training of visual foundation models [15, 16, 17, 18, 19, 20]. This progress has led to the emergence of remote sensing foundation models (RSFMs), which provide general feature representations and achieve state-of-the-art performance across various remote sensing tasks [21]. However, RSFMs still face two key challenges: (i) Compared to ImageNet-21k [22], existing RS datasets [23, 24, 25, 26] are significantly smaller (approximately 1 million vs. 14 million samples), limiting the effective MIM training of large models; and (ii) these datasets primarily capture global scene semantics [23, 24, 25] but lack the diversity and fine-grained details essential for downstream tasks, hindering the generalization of learned representations.
@@
Conventional MIM approaches, following the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [21, 31]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches (e.g., $\leq$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
Regarding the above issues, in the second part of the pipeline, we introduce an MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens (e.g., $\leq$50%) with higher HOG values for feature encoding (e.g., $\leq$15%) and pixel reconstruction (e.g., $\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence (Fig. 1). Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
@@
  1. 1.
    We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models. Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
  2. 2.
    For this pipeline, we introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training. It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
  3. 3.
    Experiment results demonstrate the effectiveness and scalability of the proposed pipeline. OpticalRS-13M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [13].
@@ 2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [26, 50], object detection [51, 52, 53], and segmentation [54, 55, 56]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. The SEN12MS dataset [57] consists of 180,662 triplets, each containing synthetic aperture radar (SAR) and multispectral Sentinel-2 image patches, and MODIS land cover maps. While SSL4EO-S12 [58] , SSL4EO-L [59] and SatlasPretrain [60] contain millions of images, the major data type in these datasets are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [61] and RS5M [62], but they primarily focus on multimodal tasks. Currently, there are also some datasets that focus solely on visible light RS images. MillionAID [23], SeCo [25] and CACo [24] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, we introduce the OpticalRS-13M dataset, which is larger and more diverse. Using this dataset for pre-training, it significantly enhances performance across multiple downstream tasks.
@@
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [78]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [27, 24, 25, 79, 80], recent advancements have primarily centered around generative self-supervised methods, especially in MIM. Specifically, many studies aim to improve generative self-supervised algorithms by leveraging general image knowledge [28, 29, 31], scaling up parameter sizes [32, 43], integrating spatio-temporal information [33, 34, 36], encompassing geometric attributes [21, 81], handling multi-sensor data [82, 83, 84, 85, 86], and employing multi-scale concepts [37, 39, 40]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RSFMs, enhancing the practicality of MIM pre-training on large-scale datasets.
Figure 3: The comparison between OpticalRS-13M and existing RS large-scale datasets in terms of data volume and image type.
@@ 3.1 Dataset Curation
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this issue, we curate a large-scale RS dataset with diverse coverage scenarios, named OpticalRS-13M. Detailed data collection and preprocessing are presented in the supplementary material.
Data Volume. OpticalRS-13M comprises 13 million high-resolution visible light RS images, establishing a significant scale advantage over existing datasets, as illustrated in Figure 3. OpticalRS-13M contains 13,203,698 images with a total of 2,630,362,174,503 pixels, averaging 199,214 pixels per image. We also offer a high-quality, lightweight version: OpticalRS-4M. OpticalRS-4M contains 3,920,829 images with a total of 688,459,799,359 pixels, averaging 175,590 pixels per image. While large-scale datasets such as SSL4EO-S12 [58] and SatLasPretrain [60] incorporate diverse data formats, e.g., including multispectral imagery, OpticalRS-13M is uniquely focused on visible light data. The dataset is carefully curated from multi-sensor acquisitions, leveraging imagery from high-resolution satellite platforms such as WorldView [87], QuickBird [88], GeoEye [89], and others. This specialization positions OpticalRS-13M as the most extensive visible light RS image repository to date, offering unparalleled utility for training vision foundation models for RS applications.
@@
Dataset Diversity. To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig. 2 (b) and (c). The analysis reveals that OpticalRS-13M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical RS datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and so on. This expanded information enhances the versatility of OpticalRS-13M for a wider range of downstream tasks.
@@ 3.2.1 Preliminaries of MAE
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder. The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error between the reconstructed targets and original patches.
@@ 3.2.3 Progressive Semantic Token Selection
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Inspired by curriculum learning  [90, 91, 92], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig. 4. In this module, we begin by selecting a limited number of patches and then select additional patches based on the similarity measure in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. $m$ is the mask ratio. We define the initial set of selected token as:
@@
Figure 6: Illustration of the patch selection process in PSTS.
@@
$\mathcal{D}(\mathbf{S}^{U},\mathbf{S}^{I})=1-\cos\left\langle\mathbf{S}^{U},\mathbf{S}^{I}\right\rangle=1-\frac{\mathbf{S}^{U}(\mathbf{S}^{I})^{\mathrm{T}}}{\lVert\mathbf{S}^{U}\rVert\cdot\lVert\mathbf{S}^{I}\rVert}$ (3)